OpenStudy (perl):

Prove that if there is an injection f : A ! A which is not surjective, then A is not a nite set. . . . . . . . . . .

OpenStudy (perl):

Prove that if there is an injection f : A -> A which is not surjective, then A is not a finite set. . . . . . . . . . .

OpenStudy (perl):

in other words, prove A is infinite

OpenStudy (perl):

i think we can do this by contradiction , assume there is an injection f A->A which is not surjective and A is finite

OpenStudy (anonymous):

Not easy lol....working on it

OpenStudy (perl):

yeah, it seems to be a stumper

OpenStudy (perl):

I got the question from this book, page 66 http://www.math.vt.edu/people/day/ProofsBook/IPaMV.pdf

OpenStudy (anonymous):

Consider the polynomials..... F(x^i)=x^(1+i) with i an element of the natural numbers....... clearly the domain which is the x^i is contained in the x^(i+1)... but of course the function is not subjective. not we know dim( Im(F(x^i)) ) + 1 = dim(x^(i+1)) now regardless of the i you choose you will always have an i+1 in the range... so thats how id argue it is not finite.

OpenStudy (anonymous):

kind of sketchy

OpenStudy (perl):

hmm

OpenStudy (perl):

what is dim x^i, and what is F(x^i) F(x^i) = x^0 + x^1 + x^2 + ...

OpenStudy (anonymous):

ok so the range is the set of polynomials x^1.... the F function acting on the domain elements F(x^i) outputs the range of polynomials of one higher order than the domain... so range is the set of polynomials x^(i+1)...... for example if i =3 then Domain = 1,x,x^2,x^3 and the range =1,x,x^2,x^3,x^4

OpenStudy (anonymous):

should be x^i not 1 in the first sentence

OpenStudy (perl):

i havent seen this function before F(x^i), is it from linear algebra

OpenStudy (anonymous):

the dimenstion of the domain of (x^i) is i +1 and the dimension of the range (x^(i+1)) would be i +2

OpenStudy (anonymous):

all the function F does is it adds 1 to the exponent

OpenStudy (anonymous):

if F is applied to x it becomes x^2....if the function F is applied to x^66 it becomes x^67

OpenStudy (perl):

ok

OpenStudy (perl):

so F(x^i) = x^i+1

OpenStudy (anonymous):

yes exactly

OpenStudy (perl):

we know dim( Im(F(x^i)) ) + 1 = dim(x^(i+1))

OpenStudy (zarkon):

prove the contrapositive

OpenStudy (perl):

the contrapositive is if A is finite then there exists a surjection

OpenStudy (zarkon):

so if \(A=\{x_1,x_2,\cdots,x_n\}\) show \(A=\{f(x_1),f(x_2),\cdots,f(x_n)\}\)

OpenStudy (perl):

errr if A is finite then for all f : A->A , f is not 1-1 or is surjective

OpenStudy (perl):

original statement: if there exists f :A->A such that f is 1-1 and not onto, then A is infinite. contrapositive: if A is finite then for all f : A->A , f is not 1-1 or is surjective

OpenStudy (perl):

do you agree with the contrapositive, did i do that right?

OpenStudy (zarkon):

not those

OpenStudy (perl):

what did i do wrong

OpenStudy (perl):

can yu finish the proof

OpenStudy (zarkon):

I guess that will work

OpenStudy (zarkon):

you can still prove it by doing what I suggested

OpenStudy (perl):

oh

OpenStudy (zarkon):

assume A is finite if f is not 1-1 then we are done so assume f is 1-1 ...prove it is surjective

OpenStudy (perl):

is that your proof that you suggested earlier ?

OpenStudy (zarkon):

no..that is just the very beginning...now do what I mentioned above

OpenStudy (zarkon):

since \(A\) is finite we can write it as \[A=\{x_1,x_2,\cdots,x_n\}\]

OpenStudy (perl):

ok

OpenStudy (zarkon):

then consider the set \[\{f(x_1),f(x_2),\cdots,f(x_n)\}\]

OpenStudy (zarkon):

can you show that each of the mappings is unique?

OpenStudy (zarkon):

given that f is one-to-one

OpenStudy (perl):

if we can show that { f(x1) ,f(x2) , ...f(xn) } = A , for any f, then we have shown it is surjective

OpenStudy (zarkon):

for any \(x_i,x_j\) with \(x_i\ne x_j\) we get \(f(x_i)\ne f(x_j)\)

OpenStudy (zarkon):

yes since each xi is mapped to a unique element in A and the mapping is exhaustive...then for any b in A there is an xi such that f(xi)=b

OpenStudy (perl):

since our original set {x1,x2,...xn} is distinct, then { f(x1),f(x2)..f(xn)} is also all distinct

OpenStudy (perl):

, because f is 1-1

OpenStudy (zarkon):

only because f is one-to-one

OpenStudy (zarkon):

yes

OpenStudy (perl):

wait, so how does that show your conclusion

OpenStudy (perl):

you said since the mapping is exhaustive

OpenStudy (zarkon):

that is the definition of onto (surjective) \[f:A\to B\] is surjective if for any \(b\in B\) there exists an \(a\in A\) such that \[f(a)=b\]

OpenStudy (zarkon):

yes \[\{f(x_1),f(x_2),\cdots,f(x_n)\}=\{x_1,x_2,\cdots,x_n\}\]

OpenStudy (zarkon):

though the order is probably different

OpenStudy (perl):

ok so we have to find , given b element B, an a element of A such that f(a) = b

OpenStudy (zarkon):

yes..that is 'onto'

OpenStudy (perl):

right, the rule might be hard to find since we could mix up the order

OpenStudy (zarkon):

yes..the function is a permutation of the elements of \(A\)

OpenStudy (zarkon):

but we don't have to specify the rule though

OpenStudy (zarkon):

though we could